The probability of hitting an aircraft is 0.001 for each shot. 

How many shots should be fired so that the probability of 

hitting with two or more shots is above 0.95?


1
Expert's answer
2022-04-12T16:47:39-0400

"P(X \ge2)=1-P(X\le1)"

"1-P(X=1)-P(X=0) \ge 0.95"

"P(X=k)=C^n_k0.001^k0.999^k"

k=0,1,2...n

So P(X=0)=0.999n and

"\mathsf P(X=1)=n \cdot 0.001\cdot0.999^{n-1}"

Thus just find the smallest n such that

"1- 0.999^{n}-n\cdot0.001\cdot 0.999^{n-1}\geq 0.95"

"0.05 \ge 0.999^{n-1}(1+0.001n)"

n=5000

"0.05 \ge 0.0067x6=0.04"